Sam purchased 500 company shares at $3.20 per share. Brokerage fees were 1.5% of the purchase price. Sam is paid a dividend of 26 cents per share, then immediately sells the shares for $4.80 each. If he pays no further brokerage fees, what is Sam’s total profit?
1
Expert's answer
2020-02-25T09:09:52-0500

Share purchase costs: 500*$3.20=$1600

Brokerage fees: $1600*1.5%=$24

Dividends on shares: $0.26*500=$130

Income from the sale of shares: 500*$4.80=$2400

Total gross profit: ($2400+$130)-($1600+$24)=$906